I'm COMPLETELY confused by this, word problems are my weak spot. :A college bookstore ordered six boxes of red pens. The store sold 32 red pens last week and 35 red pens this week. Five pens were left on the shelf. How many pens were in each box? If someone could explain how its done, i'd appreciate it
Answer by jim_thompson5910(35256)

Let x = number of pens in a single box

So each box has x pens in it.

Note: The pens are all red. So I'm just refer to them as "pens" instead of "red pens"

They ordered 6 boxes. So they have 6x pens total. Remember that 6x means "6 times x"

-------------------------------------------------------

We're told that "The store sold 32 red pens last week and 35 red pens this week".

So we know that they sold 32+35 = 67 pens so far

So they start with 6x pens and sell 67 pens. This leaves them with 6x-67 pens.

But we're also told that "Five pens were left on the shelf"

So we can say

6x - 67 = 5

since these two amounts represent what's left over.

Now solve for x


6x - 67 = 5

6x - 67+67 = 5+67 ... Add 67 to both sides

6x = 72 ... Add

6x/6 = 72/6 ... Divide both sides by 6 to isolate x

x = 12 ... Divide

So there are 12 pens in each box.


--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------